# DCD: Discriminative and Consistent Representation Distillation

### v3

@@ 4.5 Transferability of Representations
Table 4: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and Tiny ImageNet datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (Tiny ImageNet). The values in bold indicate the maximum of each row.
Table 5: Ablation results on CIFAR-100 for DCD variants with same-architecture teacher-student pairs. Two approaches are evaluated: discriminative-only loss ($\alpha=0$) and fixed temperature scaling with consistency regularization ($\alpha=0.5$, $\tau=0.07$, $b=0$).
Teacher WRN-40-2 WRN-40-2 resnet-56 resnet-110 resnet-110 resnet-32x4 VGG-13 $\Delta$
Student WRN-16-2 WRN-40-1 resnet-20 resnet-20 resnet-32 resnet-8x4 VGG-8
$\alpha=0.0$
DCD 74.37 73.28 70.52 70.58 72.63 73.80 72.44 +1.32
DCD+KD 75.51 74.24 71.57 71.39 73.62 74.86 73.78 +2.52
$\alpha=0.5$ with non-trainable $\tau=0.07$ and $b=0$
DCD 74.44 73.49 70.85 70.58 72.22 74.46 72.70 +1.63
DCD+KD 75.43 74.25 71.50 70.64 73.90 75.10 73.71 +2.45
Table 6: Ablation results on CIFAR-100 for DCD variants with cross-architecture teacher-student pairs. Two approaches are evaluated: discriminative-only loss ($\alpha=0$) and fixed temperature scaling with consistency regularization ($\alpha=0.5$, $\tau=0.07$, $b=0$).
Teacher VGG-13 ResNet-50 ResNet-50 ResNet-32x4 ResNet-32x4 WRN-40-2 $\Delta$
Student MobileNet-v2 MobileNet-v2 VGG-8 ShuffleNet-v1 ShuffleNet-v2 ShuffleNet-v1
$\alpha=0.0$
DCD 67.66 65.52 71.55 73.70 74.70 74.03 +2.46
DCD+KD 68.82 69.09 73.77 75.09 76.11 75.89 +4.40
$\alpha=0.5$ with non-trainable $\tau=0.07$ and $b=0$
DCD 68.27 66.35 72.22 74.33 75.01 74.27 +3.01
DCD+KD 69.58 69.53 74.08 74.92 75.52 76.06 +4.55
Our research focuses on transferring knowledge from a teacher network to a student network while learning representations that embody general knowledge applicable across various tasks and datasets. To investigate this, we employ a distillation process where a WRN-40-2 teacher network transfers its learned representations to a WRN-16-2 student network, which can either be trained directly from the CIFAR-100 dataset or through distillation. In our experiments, the student network functions as a fixed feature extractor, processing images from STL-10 and Tiny ImageNet, both resized to $32\times 32$. To assess the generalizability of these representations, we train a linear classifier on top of the last feature layer to perform 10-way classification for STL-10 and 200-way classification for Tiny ImageNet. The effectiveness of different distillation methods in enhancing the transferability of these representations is detailed in Table 4. Our results indicate that the DCD method, both standalone and in combination with KD, significantly outperforms all other distillation methods in improving the transferability of learned representations across different datasets. This superior transferability suggests that DCD encourages the student to learn more general and robust features, which are less overfitted to the specific training dataset and more applicable to diverse visual recognition tasks.
@@ 4.6 Ablation Study
To examine the effectiveness of consistency regularization and temperature scaling, we conduct ablation studies with two DCD variants in both Table 5 and Table 6. First, we evaluate a pure discriminative variant ($\alpha=0$) without consistency regularization. Then, we test adding consistency regularization ($\alpha=0.5$) with fixed temperature parameters ($\tau=0.07$, $b=0$). The results demonstrate that incorporating both consistency regularization and trainable temperature scaling yields the best performance, with improvements of up to 0.62% and 0.69% over the discriminative-only variant in same-architecture and cross-architecture scenarios, respectively. Additionally, removing the trainable temperature and bias parameters leads to a decrease in performance. This suggests that adaptive temperature scaling plays a crucial role in maximizing the benefits of consistency regularization during knowledge distillation. Additional ablation studies on the loss coefficients $\alpha$, $\beta$, and $\lambda$ are provided in the supplementary material.
@@ 5 Conclusions
We have presented DCD (Discriminative and Consistent Distillation), a knowledge distillation framework that combines contrastive learning with consistency regularization to enhance the traditional KD process. Our method achieves state-of-the-art performance through memory-efficient in-batch negative sampling and adaptive temperature scaling, eliminating the need for large memory banks while automatically tuning contrast levels during training. Through extensive experimentation across CIFAR-100, ImageNet, STL-10, and Tiny ImageNet datasets, we have demonstrated significant improvements over existing methods, achieving up to 20.31% relative improvement over the original KD, and 73.87% when combined with KD.

### v4

@@ 4.2 Main Results
Results on ImageNet
Table 3 presents the top-1 accuracy of student networks trained with various distillation methods on ImageNet. The results highlight the effectiveness of our approach in large-scale settings, demonstrating its ability to distill knowledge from complex models and enhance student performance. Our method consistently surpasses KD and achieves competitive results across different architectures.
Table 3: Test top-1 accuracy (%) on ImageNet validation set for student networks trained with various distillation methods across different teacher-student architectures. Results for our method are based on a single run.
Teacher Student KD [33] AT [74] SP [67] CC [58] RKD [56] CRD [66] DCD DCD+KD
[…]
Results on COCO
Table 4 shows our performance on the MS-COCO object detection task. Following [79], we adopt Faster R-CNN [60] with Feature Pyramid Network (FPN) [44] as our detection framework. We evaluate two teacher-student scenarios: ResNet-101 to ResNet-50 and ResNet-50 to MobileNet-V2. For ResNet-101 to ResNet-50 distillation, DCD+KD achieves 39.01 AP and 60.07 AP50, surpassing the baseline by +1.08 AP. When distilling from ResNet-50 to MobileNet-v2, we obtain substantial improvements of +3.81 AP over the baseline, demonstrating effective knowledge transfer even across different architectures.
Table 4: Object detection performance on MS-COCO val2017 using Faster R-CNN with FPN backbone. We use mean Average Precision (AP) and AP at IoU thresholds of 0.5 and 0.75 (AP50, AP75). Results for our method are based on a single run.
Method ResNet-101 $\rightarrow$ ResNet-18 ResNet-101 $\rightarrow$ ResNet-50 ResNet-50 $\rightarrow$ MobileNet-v2
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
Teacher 42.04 62.48 45.88 42.04 62.48 45.88 40.22 61.02 43.81
Student 33.26 53.61 35.26 37.93 58.84 41.05 29.47 48.87 30.90
KD [33] 33.97 54.66 36.62 38.35 59.41 41.71 30.13 50.28 31.35
FitNet [74] 34.13 54.16 36.71 38.76 59.62 41.80 30.20 49.80 31.69
DCD (ours) 36.98 57.44 39.79 39.01 60.07 42.88 33.28 52.97 35.15
4.3 Visualization
We present visualizations that analyze the learned representations and knowledge transfer across different distillation approaches.
Inter-class Correlations
Figure 2 evaluates the effectiveness of distillation methods on the CIFAR-100 KD task using WRN-40-2 (teacher) and WRN-40-1 (student). We compare models trained without distillation, with KL divergence [33], with CRD [66], and our proposed DCD method. DCD achieves strong correlation alignment between teacher and student logits, reducing discrepancies in their correlation matrices.
(a) Vanilla
Mean: 0.24, Max: 1.66
(b) KD [33]
Mean: 0.09, Max: 0.49
(c) CRD [66]
Mean: 0.23, Max: 1.56
(d) CRD+KD [66]
Mean: 0.10, Max: 0.57
(e) DCD (ours)
Mean: 0.22, Max: 1.59
(f) DCD+KD (ours)
Mean: 0.09, Max: 0.51
Figure 2: Correlation matrix of the average logit difference between teacher and student logits on CIFAR-100. We use WRN-40-2 as the teacher and WRN-40-1 as the student. Methods have been re-implemented according to [66].
t-SNE Visualization
Figure 3 presents t-SNE [69] visualizations of embeddings from WRN-40-2 (teacher) and WRN-40-1 (student) on CIFAR-100. Compared to standard training and [74], DCD better aligns student embeddings with the teacher, preserving structural relationships in the feature space.
(a) Teacher
(b) Vanilla
(c) KD [33]
(d) CRD [66]
(e) CRD+KD [66]
(f) DCD (ours)
(g) DCD+KD (ours)
Figure 3: t-SNE visualizations of embeddings from teacher and student networks on CIFAR-100 (first 20 classes). We use WRN-40-2 as the teacher and WRN-40-1 as the student. Methods have been re-implemented according to [66].
@@ 4.4 Ablation Study
We analyze our method through ablation experiments presented in Figures 4 and 5. We first study the effects of consistency regularization and temperature scaling, then investigate the sensitivity to hyperparameters $\alpha$, $\beta$, and $\lambda$.
(a) DCD variants without KD
(b) DCD variants with KD
Figure 4: Ablation study results on CIFAR-100. We show results for discriminative training ($\alpha=0$), discriminative and consistent training ($\alpha=0.5$, $\tau=0.07$, $b=0$), and our proposed DCD approach ($\alpha=0.5$, trainable $\tau$ and $b$). The colors correspond to each respective variant. (a) compares DCD variants without knowledge distillation, while (b) shows improvements when combined with KD. Results are based on a single run.
(a) Effect of varying $\alpha$
(b) Effect of varying $\beta$
(c) Effect of varying $\lambda$
Figure 5: Ablation study results on CIFAR-100 using WRN-40-2 as the teacher and WRN-16-2 as the student. (a) Effect of the internal DCD coefficient $\alpha$ in Equation 2. (b) Effect of DCD loss coefficient $\beta$ in Equation 7. (c) Effect of loss coefficient $\lambda$ in Equation 7. Results are averaged over five runs.
Ablation on DCD
To examine the effectiveness of consistency regularization and temperature scaling, we conduct comprehensive ablation studies on CIFAR-100, as shown in Figure 4. Starting with a pure discriminative variant ($\alpha=0$), we observe that adding consistency regularization with fixed temperature ($\alpha=0.5$, $\tau=0.07$, $b=0$) improves performance across all architectures. Our proposed method with trainable temperature parameters further enhances the results, achieving improvements of up to +1.69% without KD and +2.21% with KD over the baseline. These results demonstrate that both consistency regularization and adaptive temperature scaling contribute significantly to the method’s performance.
Ablation on $\bm{\alpha}$
We then tested different values for the loss coefficient $\alpha$ of Equation 2: $\alpha=\{0.01,0.1,0.3,0.5,0.7,1,2,5\}$ while setting $\beta=1$ and $\lambda=0$. For all the following ablations we use WRN-40-2 as teacher and WRN-16-2 as student on CIFAR-100. As shown in Figure 5(a), our method remains robust across changes in $\alpha$, with no significant difference in performance at low or high values. This robustness can be attributed to the adaptive temperature scaling mechanism, which enables automatic tuning of contrast level and logit scaling during training, providing more stable knowledge transfer.
Ablation on $\bm{\beta}$
We varied $\beta$ of Equation 7 from $\beta=0.1$ to $\beta=100$, considering values of $\beta=\{0.1,0.5,1,2,5,10,50,100\}$ while fixing $\alpha=0.5$ and $\lambda=0$. As illustrated in Figure 5(b), extremely high $\beta$ values cause significant degradation in performance due to the overwhelming contribution of the DCD loss relative to other loss terms. Very low values of $\beta$ also lead to a slight decrease in performance. The optimal range for $\beta$ is between $\beta=0.5$ and $\beta=10$, suggesting that the DCD loss should be weighted similarly to other loss terms for the best results.
Ablation on $\bm{\lambda}$
While $\lambda$ of Equation 7 is typically set to $\lambda=1.0$ [66, 7, 8], we tested values from $\lambda=0.1$ to $\lambda=10$, considering values of $\lambda=\{0.1,0.5,1,2,5,10\}$. For these experiments we fixed $\alpha=0.5$ and $\beta=1$. As shown in Figure 5(c), performance remains stable across all tested values, with the best results achieved at $\lambda=1.0$ (we also found that higher values of $\lambda=50$ and $\lambda=100$ lead to training collapse, not shown in figure). This confirms the common practice of setting $\lambda=1.0$ in prior work is optimal.
@@ 5 Conclusions
We have presented DCD, a knowledge distillation method that combines contrastive learning with consistency regularization to improve the traditional KD process. Our method achieves state-of-the-art performance through memory-efficient in-batch negative sampling and adaptive temperature scaling, eliminating the need for large memory banks while automatically tuning contrast levels during training. Through extensive experimentation across CIFAR-100, ImageNet, STL-10, and Tiny ImageNet datasets, we have demonstrated significant improvements over existing methods. Unlike previous methods such as CRD [66] that require large memory banks or WSLD [81] and IPWD [55] that focus solely on instance discrimination, our approach achieves superior performance while being more memory-efficient and capturing both local and global structural information. The effectiveness of DCD has been validated in both same-architecture and cross-architecture scenarios, with student models in several cases exceeding their teachers’ performance, particularly in WRN-40-2 to WRN-16-2 and WRN-40-2 to ShuffleNet-v1 configurations. We hope this work will inspire future research in knowledge distillation.
Acknowledgments
We would like to express our gratitude to Andreas Floros for his valuable feedback, particularly his assistance with equations and notations. We also acknowledge the computational resources and support provided by the Imperial College Research Computing Service (http://doi.org/10.14469/hpc/2232), which enabled our experiments.